Tail length in meerkats is impacted by the genotype at locus T. In a population of 2000 meerkats in the Kalahari Desert, 652 meerkats are TT, 432 are Tt, and 916 are tt. What is F (the inbreeding coefficient) at this locus?
Answer #1

Given,

TT = 652

Tt = 432 (H)

tt = 916

Total population = 2000

Now, T = 2×652 = 1304 ; T = 432, t = 432 (in Tt) ; t = 2×916 = 1832

Therefore, in 4000 population, (here T is p and t is q also p+q=1)

T = 1304+432 = 1736

So, p = 1736/4000 = 0.434

t = 1832+432 = 2264

So, q = 2264/4000 = 0.566

Now, Inbreeding coefficient (F) in a population is given by:

F = (2pq - H) / 2pq

where, H is the frequency of heterozygote in the inbred population

So, p = 0.434 and q = 0.566

2×p×q = 2×0.434×0.566 = 0.491×2000 = 982

Therefore, F = (2pq - H) / 2pq

= (982 - 432) / 982

= 0.56

​​​​​​​Hence, the inbreeding coefficient of the population is 0.56